OpenStudy (anonymous):

The equation of a curve is y=8x-x^2 Find the set of values of x for which y is greater/equal to -20.

OpenStudy (anonymous):

\[x^2 - 8x +y = 0\] As \(y \ge -20\) Take y = -20 first: \[x^2 - 8x - 20 = 0\] \[(x-10)(x+2) = 0\] \[x = 10 \quad x= -2\]

OpenStudy (anonymous):

@waterineyes mmm how do you make it in the form of inequality ? ?